# Stylos: Multi-View 3D Stylization with Single-Forward Gaussian Splatting

### v1

@@ 3.4 Style Losses
4 Experiment
Datasets. We evaluate cross-category generalization on the CO3D (Reizenstein et al., 2021) dataset by training on 17 categories and testing on 3 held-out ones, and cross-scene generalization by training on the full DL3DV-10K (Ling et al., 2024) and testing on Tanks and Temples (Knapitsch et al., 2017b). Style supervision is provided by WikiArt (WikiArt, ), with 50 images reserved for unseen-style evaluation.
Baselines. We compare with two representative open-source baselines, StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), using their released per-scene trained weights against our single DL3DV-trained model. Concurrent to Stylos, the other instant stylization solution Styl3R (Wang et al., 2025b) was proposed, but its official codes are not yet available at the time of our submission. We plan to provide comparisons once they are released.
Evaluation Metrics. Our evaluation covers three aspects. (1) To assess geometry reconstruction in the style-free setting, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) between content views and predictions, conditioned on the original first frame of each scene as the style input. (2) To measure stylization consistency, following prior work (Chiang et al., 2022), we compute LPIPS and RMSE in both short-range (consecutive frames $t$ and t1t{-}1) and long-range (distant frames $t$ and t7t{-}7) settings. (3) To further evaluate stylization quality, we report ArtScore (Chen et al., 2024), a recent metric specifically designed for reference-free evaluation of artness in generated images.
4.1 Ablation Study
Style-Content Cross Attention. As reported in Table 1, restricting cross-attention to the frame level yields only moderate performance, with a PSNR of 20.9 dB on skateboard and 19.7 dB on pizza. The hybrid variant, which combines frame-level and global attention, provides observable improvement on donut with 21.4 dB, but brings limited gains on the other categories and overall underperforms the global-only setup. By contrast, applying cross-attention at the global level consistently achieves the best results, reaching a PSNR of 21.7 dB on skateboard, 20.6 dB on pizza, and 22.1 dB on donut. Qualitative comparisons in Fig. 2 highlight these differences on the pizza example. Frame-only attention produces a poorly defined crust with noticeable artifacts, the hybrid variant recovers part of the structure but leaves the box edges blurred and unstable, while global attention delivers the most faithful result, with clear toppings and a well-preserved crust boundary.
Figure 3: Comparison of style losses on unseen donut, skateboard, and pizza scenes from the CO3D dataset. Both scene and 3D style losses yield cleaner stylized textures compared to image-level matching, while the 3D loss further conveys a stronger sense of 3D geometry.
Table 2: Consistency comparison of different style loss designs on CO3D. The 3D loss achieves the strongest short-range and long-range consistency.
Style Loss Short-range Long-range
@@ 4.1 Ablation Study
Image loss (Eq. 4) 0.046 0.035 0.147 0.135
Scene loss (Eq. 5) 0.048 0.037 0.147 0.133
3D loss (Eq. 6) 0.046 0.033 0.140 0.128
Image vs. Scene vs. 3D Style Losses. As shown in Table 2, the image-level loss achieves short-range LPIPS of 0.046 and RMSE of 0.035, but long-range consistency rises to 0.147 and 0.135. The scene-level loss slightly improves long-range RMSE to 0.133 but degrades short-range performance to 0.048 and 0.037, indicating that simple 2D aggregation reduces some long-range inconsistencies but compromises short-range accuracy. In contrast, the 3D loss applies statistics after voxel-space fusion, matching the best short-range LPIPS of 0.046, further lowering short-range RMSE to 0.033, and improving long-range scores to 0.140 and 0.128. These results indicate that enforcing style consistency in a geometry-aligned space yields sharper boundaries and more coherent textures across views. As further illustrated in Fig. 3, image-level supervision often fails to faithfully transfer style. For example, the donut lacks stylization and the wooden surface is not well synthesized, whereas scene- and 3D-level losses successfully apply the target style. Among them, the 3D style loss produces sharper boundaries and a stronger sense of 3D geometry, as observed in the skateboard and pizza examples, leading to the most coherent multi-view stylization.
From a Single View to Dozens of Views. While Stylos can process hundreds of content inputs, we qualitatively observe a gradual decrease in visual quality once the number of views per batch (denoted as “# views / batch”) exceeds 32. As shown in Fig. 4, very small batches (1 view) fail to synthesize parts of the scene such as the bench, while overly large batches (64 views) introduce edge artifacts on the tall building, potentially due to unstable Gaussian representations and the gap from training settings (no more than 24 views).
@@
StyleRF 2023 0.075 0.065 0.060 0.062 0.039 0.042
StyleGS 2024 0.033 0.038 0.031 0.034 0.038 0.037 0.069 0.061
@@
StyleRF 2023 0.116 0.110 0.176 0.154 0.096 0.092
StyleGS 2024 0.067 0.072 0.086 0.077 0.091 0.091 0.177 0.141
@@ 4.2 Comparison with State-of-the-Art
For StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), we directly evaluate with their released per-scene weights. In contrast, our Stylos is trained once on the full DL3DV dataset (Ling et al., 2024) and tested in a zero-shot manner, without ever seeing the images of these scenes. Quantitative Evaluation. According to Table 3, Stylos achieves consistently lower LPIPS and RMSE than StyleGS across both short- and long-range view pairs. This demonstrates that Stylos provides stronger cross-view consistency. Furthermore, Table 4 reports that Stylos achieves the highest ArtScore across nearly all four Tanks and Temples scenes (Knapitsch et al. (2017a)), outperforming both StyleRF and StyleGS.
@@
Method Train Truck M60 Garden
StyleRF 2023 7.71 4.12 8.48
StyleGS 2024 0.78 5.76 8.63 9.38
Stylos (ours) 9.50 9.70 9.37 9.34
Figure 5: Visual comparison between Stylos and recent per-scene 3D stylization baselines.
Qualitative Evaluation. Fig. 5 shows that Stylos successfully transfers the target style while preserving scene geometry and delivering consistent textures. In contrast, StyleGS introduces noticeable geometric distortions and texture discontinues, while StyleRF tends to over-impose style patterns, negatively impacting the object structure. Overall, Stylos achieves the best trade-off between structural fidelity and faithful stylization.
@@ 6 Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The main paper describes all model architectures in Sec. 3. We release detailed pseudo-code for all loss functions in Appendix A, and will release the full implementation upon acceptance. However, to support reproducibility, we provide detailed pseudo codes of the proposed style loss functions (image-level, scene-level, and 3D-level) in Appendix A.1. All datasets used are publicly available, as described in Sec. 4. These resources should allow researchers to reproduce and extend our findings. In addition, we provide extensive qualitative results in Appendix A.3, and analyze efficiency trends with respect to the number of views per batch in Appendix A.2. Efficiency experiments are averaged over 100 iterations per setting.
@@ Appendix A Appendix
A.1 Style Loss Pseudo Codes
We provide pseudo codes to the three variants for style losses introduced in the main paper, image-level, scene-level, and 3d style losses. This is for reproducibility purpose, and please follow the implementation in our codes which will be released in the near future.
Algorithm 1 Image-level AdaIN Style Loss
0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
0: Image-level style loss $\mathcal{L}_{\text{img}}$
1:for $i=1$ to $B$ do
2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
3:  for $j=1$ to $S$ do
4:   Extract content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,j])$
5:   Extract rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,j])$
6:   Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
7:   Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
8:   for $k=1$ to 55 do
9:    $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
10:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
11:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
12:   end for
13:  end for
14:end for
15:return $\mathcal{L}_{\text{img}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
Algorithm 2 Scene-level AdaIN Style Loss
0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
0: Scene-level style loss $\mathcal{L}_{\text{scene}}$
1:for $i=1$ to $B$ do
2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
3:  Extract batched content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,:])$
4:  Extract batched rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,:])$
5:  Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
6:  Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
7:  for $k=1$ to 55 do
8:   Reshape $F_{r}^{(k)}$ into $(B,S,C,H,W)$
9:   $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
10:   $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
11:   $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
12:  end for
13:end for
14:return $\mathcal{L}_{\text{scene}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
Algorithm 3 3D Voxel-based AdaIN Style Loss
0: Rendered features $F_{r}$, Style features $F_{s}$, Content features $F_{c}$, 3D points $P\in\mathbb{R}^{B\times S\times H\times W\times 3}$, Confidence $C$, Valid mask $M$
0: 3D style loss $\mathcal{L}_{\text{3d}}$
1:for $i=1$ to $B$ do
2:  for $k=1$ to 55 do
3:   Resize $(C,M,P)$ to match the resolution of $F_{r}^{(k)}$
4:   for $s=1$ to $S$ do
5:    $\mathcal{V}\leftarrow\text{Voxelize}(F_{r}^{(k)}[s],P[s],C[s],M[s])$
6:    $(\mu_{v},\sigma_{v})\leftarrow\text{MeanStd}(\mathcal{V})$
7:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
8:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{v}-\mu_{s}\|_{2}^{2}+\|\sigma_{v}-\sigma_{s}\|_{2}^{2}$
9:   end for
10:  end for
11:  Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
12:end for
13:return $\mathcal{L}_{\text{3d}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
@@ A.2 Efficiency Scaling
We further analyze the efficiency of Stylos by varying the number of views processed per batch. The results in Fig. 6 indicate a clear scaling pattern. For very small batches, such as 1–2 views, the average runtime per batch is low, around 0.3 s, but the measurements fluctuate heavily due to overheads and GPU scheduling. This makes single-view or extremely small-batch inference unstable, even though the raw latency is minimal.
As the batch size increases, runtime grows roughly linearly, reaching 0.75 s at 16 views and 4.83 s at 64 views. Importantly, once the batch size exceeds 4–8 views, the runtime becomes much more stable, showing only minor variation across repeated runs.
Memory usage shows a similarly consistent trend: GPU memory rises steadily from 4.16 GB at 1 view to 10.42 GB at 64 views. The growth is close to linear with respect to batch size, which makes resource requirements easy to estimate when scaling to larger workloads.
A.3 Additional Visual Results

### v2

@@ 4.1 Ablation Study
Style-Content Cross Attention. Take the pizza scene for example. Table 1 shows that the baseline method, combining frame-level and global attention, yields a PSNR of 19.78 dB and LPIPS of 0.3326. The frame variant produces slightly lower PSNR and higher LPIPS. By contrast, applying cross-attention at the global level significantly boosts the PSNR value by 0.79 dB and achieves consistently improved metrics. Qualitative comparisons in Fig. 2 highlight these differences. Frame-only attention produces a poorly defined crust with noticeable artifacts, the hybrid variant recovers part of the structure but leaves the box edges blurred, while global attention delivers the most faithful result, with clear toppings and a well-preserved crust boundary.
Figure 3: Comparison of style losses on unseen donut, skateboard, and pizza scenes from the CO3D dataset. Both scene and 3D style losses yield cleaner stylized textures compared to image-level matching, while the 3D loss further conveys a stronger sense of 3D geometry. We encourage readers to Appendix Fig. 8-10 for more visual comparisons under varing scenes and styles.
Table 2: Comparison of consistency and artistic quality among different style loss designs on CO3D. The frame stride is set to 3, and 15 held-out scenes are randomly selected for evaluation.
[…]
@@
Image loss (Eq. 3) 0.048 0.038 0.157 0.142 4.78
Scene loss (Eq. 4) 0.047 0.036 0.156 0.148 9.12
3D loss (Eq. 5) 0.047 0.034 0.153 0.142 9.15
Image vs. Scene vs. 3D Style Losses. As shown in Table 2, both scene-level and 3D style losses clearly outperform the image-level baseline in terms of artistic quality. Overall, the 3D losses exhibit superior and more stable consistency performance. These results indicate that enforcing style consistency in a voxel-level yields more coherent stylized textures across views. We observe that the image-level supervision may fail to transfer style. For example, in Fig. 3, the donut surface is not well synthesized with the expected style, whereas the scene and 3D losses successfully apply the target style. The 3D style loss produces sharper boundaries and a stronger sense of 3D geometry, as observed in the skateboard and pizza examples, leading to the most coherent multi-view stylization. We release these model weights to support reproducibility and further analysis.
From a Single View to Dozens of Views. While Stylos can process up to dozens (even hundreds) of views, we observe a gradual decrease in visual quality once the number of views per batch (denoted as “# views / batch”) exceeds 32. As shown in Fig. 4, very small batches (1 view) fail to synthesize parts of the scene like the bench, while overly large batches (64 views) introduce edge artifacts on the tall building, potentially due to the gap from our training settings (no more than 24 views). Additionally, we provide A.2 Fig. 7 to show the efficiency trend with respect to the number of views.
@@
StyleGaussian  2024 0.033 0.038 0.031 0.034 0.038 0.037 0.069 0.061
G-Style  2024 0.042 0.052 0.032 0.035 0.038 0.034 0.066 0.070
SGSST  2025 0.038 0.047 0.039 0.047 0.044 0.049 0.084 0.090
Styl3R 2025b 0.061 0.036 0.066 0.040 0.105 0.067
@@
StyleGaussian  2024 0.067 0.072 0.086 0.077 0.091 0.091 0.177 0.141
G-Style  2024 0.098 0.120 0.095 0.093 0.104 0.095 0.180 0.175
SGSST  2025 0.087 0.108 0.119 0.120 0.130 0.128 0.221 0.222
Styl3R  2025b 0.116 0.100 0.147 0.143 0.146 0.145
@@
  • StylizedGS (Zhang et al., 2025) is not included in quantitative comparisons due to its multiple failure cases observed on our test styles. Nevertheless, its quantitative results are reported in A.4 Table 5 and Table 6 for readers’ reference. Reproduced results are available at https://github.com/HanzhouLiu/Stylos.
@@ 4.2 Comparison with State-of-the-Art
To retrain StyleGaussian (Liu et al., 2024), G-Style (Kovács et al., 2024), StylizedGS (Zhang et al., 2025), and SGSST (Galerne et al., 2025), we strictly follow their released codes. Since the pretrained-weights of StyleGaussian (Liu et al., 2024) and Styl3R (Wang et al., 2025b) are publicly available, we use them to generate the visual results. Stylos is trained once on DL3DV (Ling et al., 2024), and tested in a zero-shot manner without prior knowledge of either the scenes or styles.
Quantitative Evaluation. As shown in Table 3, Stylos achieves strong and stable consistency scores, ranking the first across all consistency metrics and all four scenes. This indicates that Stylos provides markedly improved cross-view stylization consistency. Furthermore, Table 4 shows that Stylos attains either the best or second-best artistic metric values, as rated by ArtScore and ArtFID, on all four Tanks and Temples scenes (Knapitsch et al., 2017), while maintaining the fastest stylization speed. Overall, Stylos demonstrates a favorable balance between visual quality, consistency, and efficiency, suggesting its potential for practical real-time 3D stylization.
Table 4: Stylization quality, as measured by ArtScore and ArtFID (abbreviated as Score and FID respectively), and stylization time comparisons with recent 3D stylization models. Stylos achieves consistently favorable metric scores across the four scenes. Additionally, we follow StyleID (Chung et al., 2024) and calculate additional metrics as reported in A.4 Table 5 and Table 6.
@@
[…]
StyleGaussian 2024 0.78 52.79 5.76 44.93 8.63 47.48 9.38 41.14 165 m*
G-Style 2024 9.52 23.24 9.67 22.15 9.73 22.36 8.98 25.76 14.7 m*
SGSST 2025 1.84 38.24 5.34 32.34 5.26 38.73 4.89 33.54 35.2 m*
Styl3R 2025b 2.94 34.11 2.96 29.86 4.09 38.28 0.16 s
Stylos (ours) 9.50 26.40 9.70 28.71 9.37 27.44 9.34 28.06 0.05 s
    For methods with per-scene fitting required, the training phase accounts towards stylization time (see A.4).
  • As to single-forward 3D stylization approaches, training is not considered for stylization time.
Figure 5: Visual comparisons between Stylos and recent 3D stylization approaches. Stylos successfully transfers diverse artistic styles to the scenes while preserving fine structural details.
Qualitative Evaluation. Taking the truck scene with the desert-town style (yellow foreground objects with blue backgrounds) in Fig. 5 as an example, we observe that nearlly all existing approaches could reasonably produce structured results. However, their differences are discussed as follows. The zero-shot method StyleGaussian fail to generate color-consistent scenes, several portions of the truck retain undesired color blocks. Per-scene and per-style fitting approaches such as G-Style, StylizedGS, and SGSST often fail to achieve complete style transfer, the truck door remaining predominantly blue instead of adopting the target yellow hue. The other feedforward solution Styl3R similarly struggles to propagate the correct yellow color throughout the truck. In contrast, Stylos successfully renders the truck in coherent yellow while preserving clean geometry, and correctly assigns blue tones to the distant background regions, producing an appearance that closely resembles the intended desert-town aesthetic. These observations underscore the strength of Stylos in achieving both faithful style expression and robust 3D structural preservation.
4.3 Evaluation of Multi-Style Blending and Stylization Control
We validate the model’s controllable stylization capability through two interpolation experiments. First, by linearly interpolating between the embeddings of two distinct style images, we observe smooth, coherent transitions that demonstrate the model’s natural support for multi-style blending (see Fig. 6). Second, because the model is trained to reconstruct the original appearance when the style image matches the content image, we interpolate between the reconstruction (content) embedding and a stylized embedding Fig. 6. This produces a continuous spectrum of outputs with gradually increasing stylization strength, confirming that our approach enables fine-grained, post-inference control over the content–style trade-off. Together, these results show that our method supports both multi-style fusion and adjustable stylization without any additional optimization.
Figure 6: Evaluation of multi-style blending and controllable stylization. Top: interpolation between two style embeddings. Bottom: interpolation between the content embedding and a style embedding.
@@ 5 Conclusion
6 Acknowledgement
This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. The allocations on Delta and DeltaAI [allocation number CIS250529] were made by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ 7 Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The main paper describes all model architectures in Sec. 3. We provide detailed pseudo-code for the proposed 3D style loss function in Algo. 1, and will release the full implementation upon acceptance. All datasets used are publicly available, as described in Sec. 4. These resources should allow researchers to reproduce and extend our findings. In Appendix A.4, we clarify the training and inference details of the state-of-the-art methods for reproducible comparisons. In addition, we provide extensive qualitative results in Appendix A.6, and analyze efficiency trends with respect to the number of views per batch in Appendix A.2, where efficiency experiments are averaged over 100 iterations per setting.
@@ Appendix A Appendix
A.1 Implementation
Implementation Details.
For the geometry branch, we utilize VGGT as the backbone, comprising $L=24$ Alternating-Attention Transformer layers. Both the geometry transformer and the depth DPT head are initialized using pretrained VGGT weights, while all remaining layers are initialized randomly. The style aggregator similarly employs 24 Transformer layers, configured with content-global self-attention followed by content-style cross-attention. For voxelization, each output of Geo Head associates a confidence value, which is converted into a normalized intra-voxel weight via softmax over Gaussians within the same voxel; these weights determine each Gaussian’s contribution when aggregating voxel-level properties (opacity or color). The Gaussian Adapter maps the Geometry head and Style head predicted vector into explicit 3D Gaussian parameters. The module first splits the predicted vector into (i) isotropic scales, (ii) rotation quaternions, and (iii) degree-d spherical harmonic coefficients. Finally, we construct full covariance matrices using the predicted scale and rotation via the standard 3DGS covariance formulation. This yields the final set of Gaussian parameters: means, covariances, SH harmonics, opacities, scales, and normalized rotations.
In the geometry training stage, the entire network is pre-trained. The geometry outputs, poses and depth, are guided by a frozen VGGT teacher. The DINO encoder is trained to capture color and style features. For stylization fine-tuning, we optimize only the style aggregator and the style head. We set the voxel size to 0.002 for differentiable voxelization. For both geometry pre-training and stylization fine-tuning, we train the model using the AdamW optimizer for 15k iterations. We employ a cosine learning-rate schedule with a peak learning rate of $2\times 10^{-4}$ and a 1k-iteration warmup.
Data & Augmentation.
Following Anysplat (Jiang et al., 2025), we randomly sample between 2 and 20 frames per clip, maintaining a fixed total of 20 frames per GPU. The input resolution is constrained to 448 pixels on the longest side, with the aspect ratio randomized between 0.5 and 1.0. Intrinsic augmentation is applied via random center-cropping (77%–100% of the original size) and random horizontal flipping. During Geometry Pretraining, we encourage the DINO encoder to better learn color and style tokens by occasionally replacing the content image with a style image and by applying color jitter to perturb its appearance, while keeping the original image as ground truth to supervise appearance-change learning.
Loss Weights.
Style weight $\lambda_{style}=1.0$, content-reconstruction weight $\lambda_{cnt}=0.1$, CLIP consistency weight $\lambda_{clip}=1.0$, and total-variance regularization weight $\lambda_{tv}=10.0$.
@@ A.2 Efficiency Scaling
We further analyze the efficiency of Stylos by varying the number of views processed per batch. The results in Fig. 7 indicate a clear scaling pattern. For very small batches, such as 1–2 views, the average runtime per batch is low, around 0.3 s, but the measurements fluctuate heavily due to overheads and GPU scheduling. This makes single-view or extremely small-batch inference unstable, even though the raw latency is minimal. As the batch size increases, runtime grows roughly linearly, reaching 0.75 s at 16 views and 4.83 s at 64 views. Importantly, once the batch size exceeds 4–8 views, the runtime becomes much more stable, showing only minor variation across repeated runs. Memory usage shows a similarly consistent trend: GPU memory rises steadily from 4.16 GB at 1 view to 10.42 GB at 64 views. The growth is close to linear with respect to batch size, which makes resource requirements easy to estimate when scaling to larger workloads.
A.3 Ablation Analysis
Content-Style Coupling.
Stylos consists of a style pathway which contributes to the color information for the predicted 3DGS representation, and a content pathway for the other 3DGS parameters. In the style pathway, we introduce a style aggregator with cross-attention to entangle the content and style tokens. Our training pipeline consists of two stages. The first stage focuses on 3D scene reconstruction, where a random content view is used as the style reference; Stage 2 performs actual stylization using a real artistic style image. When designing Stylos, it is therefore essential to verify that introducing content-style cross-attention in the style aggregator does not harm the reconstruction process.
So, we compare three content–style coupling mechanisms (global, frame, and both) by calculating PSNR, SSIM and LPIPS between the rendered multi-view rgbs and their corresponding ground-thruth after Stage 1 training, using the first content view as the style reference. As shown in the Table 1, the global-only variant achieves the best reconstruction accuracy. This behavior is expected, because each view attends to the style tokens independently without interacting with the other views in frame attention. Treating a single view as the style reference for all views forces each view to match its patch-level representation to the same reference individually, which distorts the rendered images. Fig. 2 shows that the frame attention significantly blurs the pizza box and toppings.
Image-, Scene-, and Voxel-level Style Losses.
In this section, we provide additional visual comparisons showing that the proposed scene and 3D style losses significantly outperform the baseline image style loss. Specifically, we observe that the scene- and 3D-level losses produce cleaner and more coherent textures, such as the wooden floor in Fig. 8. As shown in Fig. 9, the 3D style loss is also able to stylize a larger portion of the table surface compared with the image loss. In Fig. 10, the seams on the tablecloth are well preserved under the 3D loss, while they become partially distorted when using the image or scene losses. We will release our implementations of all three style losses to facilitate further exploration by the community.
Figure 8: Visual comparisons of different style losses on the CO3D skateboard scene. It is clear that the 3D style loss provides a stronger sense of 3D geometry and cleaner wooden floor textures evaluated on varying style images.
Figure 9: Visual comparisons of different style losses on the CO3D pizza scene. It is clear that the 3D style loss provides stronger artistic stylization to the table surface, while the image style loss often fails to impose the expected styles to cover the whole table surface.
Figure 10: Visual comparisons of different style losses on the CO3D donut scene. It is clear that the 3D style loss provides faithful stylization to the table cover while preserving the seams on the tablecloth compared to both the image style loss and the scene style loss.
A.4 Reproducible Experiments
Reproduce State-of-the-art Models.
For StyleGaussian (Liu et al., 2024), SGSST (Galerne et al., 2025), StylizedGS (Zhang et al., 2025), and GStyle (Kovács et al., 2024), we use their officially released codebases and pretrained weights. We either run inference directly with the provided models or follow the authors’ instructions to train the models when necessary. All of them are evaluated on GH200 GPUs for fair comparisons.
We notice that StylizedGS (Zhang et al., 2025) fail in serveral style cases, producing stylized images with pure colors with barely visible geomtry details. It significantly improves its consistency scores in our experiments. For fair comparisons, we exclude these failure cases when calculating consistency and ArtScore for StylizedGS, as they could lead to very small values.
Stylization Quality Comparisons.
In Table 3 and 4, denotes that failure stylization cases for StylizedGS (Zhang et al., 2025) are excluded for fair comparisons. Please refer to our released codes for additional visual examples. Table 5 and Table 6 are provided as supplementary quantitative results to Table 4.
Table 5: Following StyleID (Chung et al., 2024), we additionally report sylization quality metrics, FID, LPIPS, LPIPS-gray, CFSD, color matching loss (HistoGAN loos), as supplementary to Table 4.
Method Train Truck
[…]
StyleGaussian 34.59 0.483 0.377 0.190 0.418 28.53 0.522 0.367 0.131 0.382
G-Style 14.80 0.471 0.364 0.165 0.185 13.65 0.512 0.353 0.087 0.176
StylizedGS 22.90 0.707 0.648 0.216 0.396 22.65 0.779 0.713 0.084 0.417
SGSST 24.15 0.520 0.409 0.220 0.257 19.50 0.577 0.445 0.177 0.196
Styl3R 19.77 0.670 0.598 0.244 0.364 19.28 0.682 0.575 0.102 0.350
Stylos (ours) 15.30 0.620 0.529 0.223 0.241 16.67 0.625 0.471 0.084 0.237
Table 6: Following StyleID (Chung et al., 2024), we additionally report sylization quality metrics, FID, LPIPS, LPIPS-gray, CFSD, color matching loss (HistoGAN loos), as supplementary to Table 4.
Method M60 Garden
[…]
StyleGaussian 30.54 0.506 0.413 0.138 0.467 25.23 0.569 0.454 0.189 0.480
G-Style 13.93 0.498 0.395 0.092 0.208 16.17 0.500 0.364 0.103 0.179
StylizedGS 28.33 0.815 0.728 0.102 0.443 33.73 0.876 0.827 0.074 0.570
SGSST 23.95 0.552 0.458 0.204 0.264 20.93 0.529 0.415 0.233 0.228
Styl3R 17.14 0.646 0.573 0.124 0.314 22.38 0.637 0.556 0.097 0.335
Stylos (ours) 16.61 0.558 0.457 0.098 0.252 16.26 0.625 0.509 0.080 0.242
Stylization Time Comparisons.
We run 48 samples as a hardware warm-up phase, and calculate the average inference speeds over the following 112 samples. For fair comparisons, all methods are evaluated on the same hardware, using an NVIDIA GH200 GPU with identical configurations. The reported stylization time of the per-scene fitting required method considers per-scene training time besides the rendering time. We strictly follow the original authors’ training setups and inference pipelines. All 3DGS-based methods (Liu et al., 2024; Galerne et al., 2025; Zhang et al., 2025; Kovács et al., 2024) are evaluated on the full-resolution images, whereas single-forward methods, including Styl3R (Wang et al., 2025b) and Stylos, are evaluated at their preset input resolutions, i.e., $256\times 256$ and $448\times 448$, respectively.
(a) The reconstructed metal bars on the gate are unclear as shown in the mid row, and the stylized metal bars at the top row are barely visible either.
(b) The green colors fail to correspond to the intended regions in the reconstructed scene (at the mid row), and the stylized multi-view images mistakenly interpret the green area as part of the target pattern, leading the model to apply stylization to it, (at the top row).
Figure 11: Failures caused by inaccurate scene reconstruction, such as missing geometric details or incorrect region colors. Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
(a) The reconstructed scene is generally accurate. However, our model introduces a pure white region in the bottom-left area (at the top row), which does not appear in the original style image.
(b) The reconstructed scene is generally accurate. However, our model hallucinates a pure white region on the left side (at the top row), which is absent from the original style image.
Figure 12: Failures caused by generating colors that are absent from the original style images. Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
(a) Our model produces a blurred stylized region in the upper-left area of the results (as shown in the top row).
Figure 13: Failures caused by introducing blurry or over-smoothed style colors. Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
A.5 Limitations
Stylos tends to underperform on scenes dominated by high-frequency or highly cluttered structures, such as dense foliage, thin branches, or wire-like elements, where preserving fine geometric detail is particularly challenging. The method also struggles with style categories that introduce strong global lighting shifts or extreme color palettes, which can lead to over-saturation or loss of subtle appearance cues. In addition, performance decreases as the number of input views increases. This degradation primarily stems from limitations of the VGGT backbone: as view count grows, its underlying geometric reconstruction becomes less stable, which in turn negatively affects the quality of the stylized outputs. More specifically, the failure cases can be grouped into three categories: (1) Failures caused by inaccurate scene reconstruction, such as missing geometric details or incorrect region colors; (2) Failures caused by generating colors that are absent from the original style images. (3) Failures caused by introducing blurry or over-smoothed style colors.
Figure 14: Style images from WikiArt (2025) used for model evaluation in all quantitative comparisons, while not used for training Stylos.
Figure 15: Style images from WikiArt (2025) used for model evaluation in all quantitative comparisons, while not used for training Stylos.
A.6 Additional Visual Results